a radioactive isotope has a half life of 6 yrs. a sample of this isotope has a mass of 12mg. what’s the mass of the isotope after 18 yrs?
1
Expert's answer
2020-12-17T04:50:57-0500

"N=N_0e^{-\lambda t}\to m=m_0e^{-\lambda t}\to m=m_0e^{-\ln2\cdot t/T_{1/2}}\to"


"m=12e^{-\ln2\cdot 18/6}=1.51(mg)"